Assume that Seminole, Inc., considers issuing a Singapore dollar‑denominated bond at its present coupon rate of 7 percent, even though it has no incoming cash flows to cover the bond payments. It is attracted to the low financing rate, since U. S. dollar-denominated bonds issued in the United States would have a coupon rate of 12 percent. Assume that either type of bond would have a four­‑year maturity and could be issued at par value. Seminole needs to borrow $10 million. Therefore, it will either issue U. S. dollar denominated bonds with a par value of $10 million or bonds denominated in Singapore dollars with a par value of S$20 million. The spot rate of the Singapore dollar is $.50. Seminole has forecasted the Singapore dollar’s value at the end of each of the next four years, when coupon payments are to be paid:

                        End of Year                 Exchange Rate of Singapore Dollar

                                 1                                       $.52

                                 2                                        .56

                                 3                                        .58

                                 4                                        .53

Determine the expected annual cost of financing with Singapore dollars. Should Seminole, Inc., issue bonds denominated in U.S. dollars or Singapore dollars? Explain (Using IRR to find the %).
Answer #1

The bond amount of $ 10 million if raised in US dollars would cost 12% (as the bond can be issued at par)

However, if the same amount is to be raised in Singapore dollar, at an exchange rate of S$1 = $0.5 S$20 million ( = $ 10 millin) bonds needs to be issued at 7%.

After end of each year (1-4) an amount of S$ 20 milion * 7% = S$1.4 million

And at end of 4th year an amount of S$20 million as redemption value needs to be paid.

As there is no source of cash inflows to cover the bond payments . So, amount in US dollars need to be converted to Singapore dollars at the end of each year for payment of coupon and at the end of 4th year for redemption value

Amount of US dollars required at end of year1 = S$1.4 million * 0.52 = $0.728 million

Amount of US dollars required at end of year2 = S$1.4 million * 0.56 = $0.784 million

Amount of US dollars required at end of year3 = S$1.4 million * 0.58 = $0.812 million

Amount of US dollars required at end of year4 = S$1.4 million * 0.53 = $0.742 million

Amount of US dollars required at end of year4 = S$20 million * 0.53 = $10.6 million (for redemption amount)

So, for the bond issued in Singapore dollars, the cost r by the Bond pricing formula is (all figures in $ million)

10 = 0.728/(1+r) + 0.74 / (1+r)2 + 0.812/(1+r)3 + (0.742+10.6)/(1+r)4

By hit and trial

Taking r= 10%

RHS = 0.728/(1+0.1) + 0.74 / (1+0.1)2 + 0.812/(1+0.1)3 + (0.742+10.6)/(1+0.1)4

= 9.66 which is less than the LHS

Taking r= 9%

RHS = 0.728/(1+0.09) + 0.74 / (1+0.09)2 + 0.812/(1+0.09)3 + (0.742+10.6)/(1+0.09)4

= 9.9897 which is very slighty less than the LHS

Taking r= 8.9%

RHS = 0.728/(1+0.089) + 0.74 / (1+0.089)2 + 0.812/(1+0.089)3 + (0.742+10.6)/(1+0.089)4

= 10.02284 which is slighty higher than the LHS

So, the annual cost of financing with Singapore dollars is (by Linear approximation method)

r = 8.9% + (10.02284- 10)/ (10.02284 - 9.9897) * (9%- 8.9%)

= 8.96892 %

=8.97%

Since this cost is less than the cost of financing bond in US dollars (12%). Seminole Inc should issue bonds in Singapore Dollars